OpenStudy (anonymous):

i need help with this question Given that f(x) = 5x − 10 and g(x) = x + 3, solve for f(g(x)) when x = −1

OpenStudy (xapproachesinfinity):

\(f(x)=5x-10\) then \(f(\square )=5\square-11\) \(f(\bigstar)=5\bigstar-11\) so \(f(\color{red}{g(x)})=5?-11\)

OpenStudy (xapproachesinfinity):

what should i put in place of question mark? doing the same thing i did with the star, with the square or even x as it was already

OpenStudy (anonymous):

im not totally sure because this is exactly wat it asking me Given that f(x) = 5x − 10 and g(x) = x + 3, solve for f(g(x)) when x = −1. −30 0 10 12

OpenStudy (xapproachesinfinity):

first thing find \(f(g(x))\) following what i given as examples okay we will go to when x=1 after

OpenStudy (anonymous):

oh ok then so would it possibly be maybe 0 ?

OpenStudy (anonymous):

or 10 ?

OpenStudy (xapproachesinfinity):

he forgot the given option okay read my reply above and look at it and see if you understand what you should do

OpenStudy (xapproachesinfinity):

let me ask you a question evaluate f(2)? how would you do that

OpenStudy (anonymous):

ummm i think you woukd square ?

OpenStudy (xapproachesinfinity):

square what? our function is \(f(x)=5x-10\) there nothing to square here pay attention \(f(\color{magenta}{x})=5\color{magenta}{x}-10\) \(f(\color{magenta}{2})=5(\color{magenta}{2})-10\) what i did here is replace x with 2

OpenStudy (xapproachesinfinity):

so what is \(f(\color{red}{2+t})\)

OpenStudy (xapproachesinfinity):

we are trying to learn this so pay some attention okay^_^

OpenStudy (anonymous):

ohhhh ok so then it mostly multiplying right ? to wher it could be possible 10 or 12 ?? and lol im sorry im paying attention carefullly

OpenStudy (anonymous):

im sorry im not good with alegebra 2 like at all have to a exam 34 questions and this is one of them online

OpenStudy (xapproachesinfinity):

hey you don't seem to pay attention i said forgot about the options you were given! we are trying to build this up and eventually we will get to the answer

OpenStudy (anonymous):

hey im trying and im full on attention . i never understand this type of stuff

OpenStudy (xapproachesinfinity):

that's valid! there is no such thing as i'm not good at something we all start learning at some point! the more you struggle with something, the more you learn it!

OpenStudy (anonymous):

true buttttt im not great doing alegebra 2 online on virtual school like legit man !!!

OpenStudy (xapproachesinfinity):

what i'm saying is put instead of x anything else that they want you to evaluate like a number (can be any number) or expression like \(2+t\)

OpenStudy (anonymous):

oh ok i think

OpenStudy (xapproachesinfinity):

hey just forgot for the moment that you have troubles with algebra empty you head with those thoughts! make your self ready to learn! doesn't matter for me if you are good at this or not okay! our focus is solve this task^_^

OpenStudy (anonymous):

lolz okay dokie ^_^

OpenStudy (xapproachesinfinity):

Now! follow my function is \(f(x)=5x-10\) this for x (x means for any number) i want to find f(2), now 2 is a specific input all i need is to plug in 2 (replace x by 2 anywhere i see x i will put 2) so i got \(f(2)=5\times2-10\)

OpenStudy (xapproachesinfinity):

i can do this for any other number and also an expression like 2+w this an expression

OpenStudy (xapproachesinfinity):

you understand so far?

OpenStudy (anonymous):

a little bit

OpenStudy (xapproachesinfinity):

okay so what is \(f(3)\)

OpenStudy (anonymous):

3

OpenStudy (xapproachesinfinity):

let's go slowly then hehe

OpenStudy (xapproachesinfinity):

look what i did for 2 do the same thing okay

OpenStudy (xapproachesinfinity):

pause and put some thinking about it okay! pls try to motivate yourself to learn this! it is not hard as you look at it it is you that convinced yourself to believe that this is hard

OpenStudy (anonymous):

grrrrr ok lol

OpenStudy (xapproachesinfinity):

okay take a good look at \(f(\color{blueviolet}{2})=5\times\color{blueviolet}{2}-10\) you need to do the same thing with 3

OpenStudy (anonymous):

ok

OpenStudy (xapproachesinfinity):

show me your work okay

OpenStudy (anonymous):

i give up lol legit im sorry but i do thanks anyways but im just gionna take another guess again grrr >_<

OpenStudy (xapproachesinfinity):

hehehe let me tell you a short story! when i started learning math in particular algebra my teacher used to ask me what is \(5^2\) i responded sometimes 10 sometimes 15. i was really bad at the subject! but i never give up! i struggled to learn it. by the end of the semester I turned my worst subject to a subject that i get interested in more and more!

OpenStudy (xapproachesinfinity):

I'm still bad! so don't ever look down. we are made to learn and progress so don't be lazy! try! put your efforts on it!

OpenStudy (xapproachesinfinity):

my advice to you is start thinking about things in a different positive way^_^